# MG-SLAM: Structure Gaussian SLAM with Manhattan World Hypothesis

### v2

@@ Abstract
Gaussian SLAM systems have made significant advancements in improving the efficiency and fidelity of real-time reconstructions. However, these systems often encounter incomplete reconstructions in complex indoor environments, characterized by substantial holes due to unobserved geometry caused by obstacles or limited view angles. To address this challenge, we present Manhattan Gaussian SLAM (MG-SLAM), an RGB-D system that leverages the Manhattan World hypothesis to enhance geometric accuracy and completeness. By seamlessly integrating fused line segments derived from structured scenes, MG-SLAM ensures robust tracking in textureless indoor areas. Moreover, The extracted lines and planar surface assumption allow strategic interpolation of new Gaussians in regions of missing geometry, enabling efficient scene completion. Extensive experiments conducted on both synthetic and real-world scenes demonstrate that these advancements enable our method to achieve state-of-the-art performance, marking a substantial improvement in the capabilities of Gaussian SLAM systems.
Figure 1: MG-SLAM leverages line segments to achieve SOTA results in camera pose estimation and scene reconstruction. Additionally, by applying structural surface constraints, we enhance and complete the scene through the interpolation of new Gaussians for absent geometry.
1 Introduction
Simultaneous Localization and Mapping (SLAM) is a fundamental problem in computer vision that aims to map an environment while simultaneously tracking the camera pose. Learning-based dense SLAM methods, particularly neural radiance field (NeRF) approaches [51, 22, 41, 18, 52, 36, 8], have demonstrated remarkable improvements in capturing dense photometric information and providing accurate global reconstruction over traditional systems based on sparse point clouds or voxels [33, 35, 32, 14, 9, 17]. However, NeRF methods still face drawbacks such as over-smoothing, bounded scene representation, and computational inefficiencies. Recently, Gaussian-based SLAM [19, 28, 46, 26, 10, 25] has emerged as a promising approach utilizing volumetric radiance fields. Leveraging explicit 3D Gaussian representation [20], Gaussian SLAMs deliver high-fidelity rendering and fine-grained scene reconstruction, overcoming the limitations of NeRF-based methods.
@@ 1 Introduction
To overcome these challenges, we leverage the renowned Manhattan World hypothesis [5] as a foundational strategy for refining and completing scene geometries. This assumption posits that the built environment predominantly adheres to a grid-like structure, with surfaces and lines aligning with three orthogonal directions. These lines and planar surfaces impose meaningful constraints on the tracking and mapping processes in the Gaussian SLAM system.
Specifically, we encompass enhancements in tracking, mapping, and scene completion. In tracking, we utilize line features derived from the structured scenes as robust feature foundations in textureless areas, backprojecting and reprojecting these line segments for pose optimization and full bundle adjustment. In mapping, we apply a photometric loss for the reprojected line features to refine the map. This approach ensures that the reconstructed scene adheres closely to the true structure of the environment, thereby improving both its geometry accuracy and rendering quality.
Furthermore, the Manhattan World hypothesis [5] facilitates the identification and interpolation of structured surfaces, such as floors and ceilings. These planar surfaces are critical to defining the overall geometry of space but are often partially obscured or missing from the captured views. By segmenting these incomplete surfaces—refined by the extracted lines as boundaries—we can predict their continuation beyond the directly observed portions by generating new Gaussians. This strategy enables us to optimize the representation of large surfaces within the scene, enhancing the completeness of the rendered map. Finally, we compress the Gaussian representation into mesh surfaces by incorporating regularization terms through Poisson reconstruction [15]. This approach enables the extraction of high-quality mesh, previously unavailable in the Gaussian SLAM systems, making it readily available for downstream tasks.
Overall, our work presents several key contributions, summarized as follows:
  • We propose MG-SLAM, a novel RGB-D Gaussian SLAM system that capitalizes on the Manhattan World hypothesis [5]. This assumption introduces lines and planar surfaces for robust tracking, map refinement, and surface completion for neural-dense SLAM systems.
@@
  • Extensive experiments conducted on both large-scale synthetic and real-world datasets demonstrate that our system offers state-of-the-art (SOTA) tracking and comprehensive map reconstruction, achieving 50% lower ATE and 5dB enhancement in PSNR on real-world scenes, meanwhile operating at an exceptionally high frame rate. These advancements remarkably outperform previous Gaussian SLAM systems.
2 Related Work
2.1 Neural dense SLAM
@@
Neural Implicit SLAM systems leveraging NeRF [30] are adept at handling complex topological structures by employing differentiable scene representation approaches [51, 4, 24, 18, 36, 11, 49, 12]. Despite these advancements, NeRF methods often struggle with the over-smoothing issue, where fine-grained object-level geometry and features are difficult to capture during reconstruction [26]. Moreover, these methods suffer from catastrophic loss as the scenes are implicitly represented by MLPs.
@@
In contrast, the high-fidelity and fast rasterization capabilities of 3D Gaussian Splatting [20] enable higher quality and efficiency on scene reconstruction [44, 26, 39]. MonoGS [28] utilizes a map-centric SLAM approach that employs 3D Gaussian representation for dynamic and high-fidelity scene capture. SplaTAM [19] adopts an explicit volumetric approach using isotropic Gaussians, enabling precise camera tracking and map densification. However, existing Gaussian SLAM systems lack effective camera pose optimization, limiting tracking accuracy in complex scenes. Moreover, they struggle to effectively map unobserved areas, often resulting in incomplete reconstructions with gaps and holes. This limitation becomes more problematic in settings where the camera’s movement is restricted, leading to significant unmodeled areas in structured indoor scenes. Additionally, current Gaussian-based SLAM methods are unable to generate meshes due to the discrete nature of 3D Gaussians which poses difficulties in surface extraction.
2.2 Reconstruction with structure optimization
Line features are known to remarkably enhance camera pose optimization [50, 43]. Traditional SLAM systems [42, 14, 48, 3] utilize the point feature and line segment for capturing high-level geometric elements and structural properties, which accurately refine the reconstructed map. However, these systems rely on sparse map expressions and struggle to interpret dense surfaces. The integration of line features in neural dense SLAM systems still remains an unexplored topic.
As for surface interpolation, several offline reconstruction approaches have been proposed. [16] introduced a NeRF-based method that optimizes reconstructed scenes by aligning the signed distance field (SDF) of planar surfaces. [40] and [27] proposed methods for geometric-aware Gaussian initialization from point clouds, which effectively preserve structural correlations, especially in textureless areas. However, these approaches are fundamentally based on optimization from observed views, which limits their ability to interpolate unseen structured geometry in reconstructed scenes.
3 Method
Figure 2 illustrates the pipeline of our proposed method. Under the constraints of the Manhattan World hypothesis, MG-SLAM introduces line segments and structured surfaces to enhance camera pose estimation and map reconstruction. Section 3.2 details the tracking mechanism that utilizes both point and line features. We utilize a specific strategy for fusing line segments to ensure reliable identification of line features. Section 3.3 discusses the Gaussian representation, including a specialized loss term dedicated to the reconstruction of line segments. Section 3.4 and section B.2 describe the completion and refinement of the scene, grounded in the assumption of structured surfaces. Section B.3 describes the mesh generation utilizing regularization losses.
@@ 3 Method
Figure 2: The two-phase pipeline illustration of the proposed MG-SLAM. The upper section visualizes the parallel processes of the tracking and mapping systems. The lower section presents the post-optimization of scene completion and mesh extraction.
3.1 Notation
@@
3.2 Tracking
We utilize the backprojection of point features and line segments extracted from 2D images into 3D space in parallel for optimizing the camera pose, based on the approach outlined in PLVS [14]. Moreover, we incorporate strategies for line segment fusion and suppression, designed to merge shorter segments into longer ones and eliminate unstable lines. This method is grounded in the understanding that longer line features tend to offer greater reliability and robustness throughout the tracking process.
3.2.1 Point reprojection error
@@
3.2.2 Line reprojection and backprojection error
@@
$\mathcal{L}_{\rm 2D}^{l}=\begin{bmatrix}\left\|d_{k}\cdot\pi(R_{k}\mathcal{X}_{p}^{w}+t)\right\|\\ \left\|d_{k}\cdot\pi(R_{k}\mathcal{X}_{q}^{w}+t)\right\|\end{bmatrix}_{\Sigma_{\rm 2D}}\in\mathbb{R}^{2}~{},~{}~{}d_{k}=\frac{\bar{\mathcal{P}_{p}}\times\bar{\mathcal{P}_{q}}}{\left\|\bar{\mathcal{P}_{p}}\times\bar{\mathcal{P}_{q}}\right\|}|_{k}$ (2)
@@
which considers the cross-product of the differences between the point’s position and the back-projected positions of the line’s endpoints, normalized by the distance between these back-projected endpoints. Additionally, the distance between the world domain point $\mathcal{X}^{w}_{i}$ and its corresponding back-projected point in the image plane $\mathcal{P}_{i}\in\mathbb{R}^{2}$ is determined by:
@@
3.2.3 Line segment fusion
Figure 3: The ablation of the line segment extraction outcomes for our method on the scene frl_apartment_4 from the Replica Apartment dataset [38].
The stability of tracking can be notably enhanced by the presence of elongated line segments. However, the extraction process, particularly through EDLines, often yields fragmented segments as a consequence of various disturbances like image noise. Our method integrates the following key steps: (1) Fusing line segments that are directionally aligned within a one-degree angle difference, ensuring they follow the same path. (2) Ensuring these segments are close—within a 10-pixel distance of each other’s nearest endpoints—yet do not overlap, preserving their distinctness while allowing for precise merging. (3) Verifying the vertical distance between the corresponding endpoints of one segment to the entirety of another is less than certain pixel threshold, maintaining geometric consistency. Only segments fulfilling all three criteria are merged, producing longer and more reliable lines. Furthermore, we filter out segments that fall below a predefined length threshold, relative to the image size, to improve the system’s tracking robustness. Figure 3 displays the ablation results for our feature extraction approach for 3D line segments, highlighting the effectiveness of the backprojection, line fusion, and filtering strategies.
3.2.4 Full bundle adjustment error
The overall objective function minimized during the full bundle adjustment is given by:
$\mathcal{L}_{\rm BA}=\sum_{\mathcal{K}}\sum_{\mathcal{U}}\rho\left(\left\|\mathcal{L}_{\rm 2D}^{p}\right\|_{\Sigma_{\mathcal{P}}}\right)+\sum_{\mathcal{V}}\rho\left(\left\|\mathcal{L}_{\rm 2D}^{l}\right\|_{\Sigma_{\rm 2D}}\right)+\sum_{\mathcal{V}}\rho\left(\left\|\mathcal{L}_{\rm 3D}^{l}\right\|_{\Sigma_{\rm 3D}}\right)$ (6)
Here, $\mathcal{K}$ represents the set of chosen keyframes, while $\mathcal{U}$ and $\mathcal{V}$ denote the sets of points and lines extracted within the current frame. The Huber cost function $\rho(\left\|e\right\|)_{\Sigma^{-1}})=e^{T}\Sigma^{-1}e$ is applied to mitigate the influence of outliers. The optimization process utilizes the Levenberg-Marquardt method, which solves the augmented normal equations by iteratively updating the parameters $\Theta$ as:
$(J^{T}\Sigma^{-1}J+\lambda\mathbb{I})\Delta\Theta=-J^{T}\Sigma^{-1}\epsilon$ (7)
where $\epsilon$ aggregates the reprojection and backprojection errors defined in Equation 6, and $\Sigma$ combines the covariance matrices relevant to these errors.
3.3 Mapping
The map representation is based on Gaussians ellipsoids $\mathcal{G}=\alpha\mathcal{N}(\mu^{w},\Sigma^{w})$, where $\alpha\in[0,1]$ is the opacity, $\mu^{w}$ and $\Sigma^{w}$ are mean and covariance in world coordinate. Each $\mathcal{G}$ is associated with the color of appearance $\mathcal{C}$ and semantic color $\mathcal{S}$. Semantic segmentation is utilized to identify the surface such as floors for structure optimization and surface extraction, explained in Section 3.4.
3.3.1 Scene representation

### v3

# MG-SLAM: Structure Gaussian Splatting SLAM with Manhattan World Hypothesis

@@ Abstract
Gaussian Splatting SLAMs have made significant advancements in improving the efficiency and fidelity of real-time reconstructions. However, these systems often encounter incomplete reconstructions in complex indoor environments, characterized by substantial holes due to unobserved geometry caused by obstacles or limited view angles. To address this challenge, we present Manhattan Gaussian SLAM, an RGB-D system that leverages the Manhattan World hypothesis to enhance geometric accuracy and completeness. By seamlessly integrating fused line segments derived from structured scenes, our method ensures robust tracking in textureless indoor areas. Moreover, The extracted lines and planar surface assumption allow strategic interpolation of new Gaussians in regions of missing geometry, enabling efficient scene completion. Extensive experiments conducted on both synthetic and real-world scenes demonstrate that these advancements enable our method to achieve state-of-the-art performance, marking a substantial improvement in the capabilities of Gaussian SLAM systems.
Note to Practitioners—This paper was motivated by the limitations of Gaussian Splatting SLAM systems in complex indoor environments, where textureless surfaces and obstructed views often lead to substantial tracking errors and incomplete maps. While existing systems excel in high-fidelity reconstruction, they struggle with frame-to-frame or point-feature tracking in large-scale environments, particularly with significant camera rotations and obscured structures. In this paper, we enhance the neural dense SLAM by integrating the Manhattan World hypothesis, applying its parallel line and planar surface constraints for more robust tracking and mapping. We incorporate line segment features into both tracking and mapping to improve structural accuracy. Moreover, we propose a post-optimization method that interpolates new Gaussian primitives to effectively fill gaps on planar surfaces. Extensive experiments on multiple datasets demonstrate the superiority of our approach in large-scale indoor environments, resulting in more accurate tracking and mapping.
Index Terms:
SLAM, 3D Gaussian Splatting, Manhattan World.
I Introduction
Simultaneous Localization and Mapping (SLAM) is a fundamental problem in computer vision that aims to map an environment while simultaneously tracking the camera pose. Learning-based dense SLAM methods, particularly neural radiance field (NeRF) approaches [1, 2, 3, 4, 5, 6, 7, 8], have demonstrated remarkable improvements in capturing dense photometric information and providing accurate global reconstruction over traditional systems based on sparse point clouds [9, 10, 11, 12, 13, 14, 15]. However, NeRF methods still face drawbacks such as over-smoothing, bounded scene representation, and computational inefficiencies [16, 17, 18, 19]. Recently, Gaussian-based SLAM [20, 21, 22, 23, 24, 25] has emerged as a promising approach utilizing volumetric Gaussian primitives [26]. Leveraging these explicit representations, Gaussian SLAMs deliver high-fidelity rendering and fine-grained scene reconstruction, overcoming the limitations of NeRF-based methods.
Figure 1: Visualization of MG-SLAM on scene0000_00 and scene0207_00 of the ScanNet dataset [27]. Our method leverages robust line segments to achieve superior camera pose estimation and scene reconstruction results. Moreover, by applying structural surface constraints, we enhance and complete the planar surfaces of the scene through the insertion of new Gaussian primitives to fill gaps.
@@ I Introduction
To overcome these challenges, we leverage the renowned Manhattan World (MW) hypothesis [28] as a foundational strategy for refining and completing scene geometries. This assumption posits that the built environment predominantly adheres to a grid-like structure, with surfaces and lines aligning with three orthogonal directions. These lines and planar surfaces impose meaningful constraints on the tracking and mapping processes in the Gaussian SLAM system.
Specifically, we encompass enhancements in tracking, mapping, and planar surface completion. In tracking, we utilize fused line features derived from the structured scenes as robust feature foundations in textureless areas, backprojecting and reprojecting these line segments for pose optimization and full bundle adjustment. In mapping, we apply a photometric loss for the reprojected line features to refine the map. This approach ensures that the reconstructed scene adheres closely to the true structure of the environment, thereby improving both its geometry accuracy and rendering quality.
Furthermore, the MW hypothesis facilitates the identification and interpolation of structured surfaces, such as floors and ceilings. These planar surfaces are critical to defining the overall geometry of indoor scenes but are often partially obscured or missing from the captured views [29, 30]. By segmenting these incomplete surfaces—refined by the extracted lines as boundaries—we can predict their continuation beyond the directly observed portions by generating new Gaussians. This strategy enables us to optimize the representation of large surfaces within the scene, enhancing the completeness of the rendered map. Finally, we compress the Gaussian representation into mesh surfaces by incorporating regularization terms through Poisson reconstruction [31]. This approach enables the extraction of high-quality mesh, previously unavailable in the Gaussian SLAM systems, making it readily available for downstream tasks.
Overall, our work presents the following key contributions:
  • We propose MG-SLAM, a novel RGB-D Gaussian SLAM system that capitalizes on the MW hypothesis. This assumption introduces lines and planar surfaces for robust tracking, map refinement, and surface completion for neural-dense SLAM systems.
  • We incorporate line segments along with an additional fusion and filtering strategy into the neural-dense SLAM system, effectively improving its tracking capabilities in textureless indoor environments and enhancing the quality of the dense Gaussian map.
@@
  • Extensive experiments conducted on both large-scale synthetic and real-world datasets demonstrate that our system offers state-of-the-art (SOTA) tracking and comprehensive map reconstruction, achieving 50% lower ATE and 5dB enhancement in PSNR on large-scale Apartment dataset, meanwhile operating at a high frame rate.
II Related Work
II-A Neural Dense SLAM
@@
Neural Implicit SLAM systems [1, 32, 4, 3, 6, 33, 34, 35, 36] leveraging NeRF [37] are adept at handling complex reconstruction using implicit volumetric representation. Despite these advancements, NeRF methods often struggle with the over-smoothing issue, where fine-grained object-level geometry and features are difficult to capture during reconstruction [23]. Moreover, these methods suffer from catastrophic loss as the scenes are implicitly represented by MLPs.
@@
In contrast, the high-fidelity and fast rasterization capabilities of 3D Gaussian Splatting [26] enable higher quality and efficiency on scene reconstruction [20, 38, 21, 24, 39, 23, 40, 41, 42, 43, 44]. MonoGS [21] utilizes a map-centric SLAM approach that employs 3D Gaussian representation for dynamic and high-fidelity scene capture. SplaTAM [20] adopts an explicit volumetric approach using isotropic Gaussians, enabling precise camera tracking and map densification. Photo-SLAM [24] and RTG-SLAM [39] integrate the traditional feature-based tracking system [45, 46] with Gaussian mapping, providing robust tracking and exceptional real-time processing capabilities. However, existing Gaussian SLAM systems lack effective camera pose optimization in textureless environments, limiting tracking accuracy in indoor scenes. Moreover, they struggle to effectively reconstruct unobserved areas, often resulting in incomplete reconstructions with gaps and holes. This limitation becomes more problematic in settings where the camera’s movement is restricted, leading to significant unmodeled areas in structured indoor scenes. Additionally, current Gaussian-based SLAM approaches face challenges in direct mesh generation due to the discrete nature of 3D Gaussian primitives, which complicates surface extraction. To address this, recent Gaussian-based systems [38, 22, 42] apply TSDF fusion [47] on rendered images to produce meshes. However, this approach is limited to observed viewpoints, leaving unobserved regions incomplete. More critically, it fundamentally depends on the offline volumetric TSDF projection using rendered images, which is independent of the reconstructed Gaussian maps. To overcome these challenges, our method incorporates line segments and planar surface assumption to seamlessly fill the gaps of structured surfaces and directly extract high-quality mesh from volumetric Gaussian representations.
II-B SLAM with Structure Optimization
Line features are known to significantly enhance camera pose optimization by capturing high-level geometric elements and structural properties [48, 49]. Traditional SLAM systems [50, 51, 13, 52, 53, 54, 55, 56] combine point features, line segments, or planar surfaces to refine camera pose estimation and improve map reconstruction. These approaches are particularly effective in dynamic or textureless environments, where point-based methods often face substantial challenges.
Building on the MW hypothesis, recent advancements in SLAM systems leveraging planar constraints have further mitigated tracking and mapping drift. For instance, [57] demonstrates accurate camera pose estimation and sparse 3D map generation in monocular SLAM. [58] achieves drift-free rotational motion estimation by leveraging structural regularities captured by line features. ManhattanSLAM [59] extends this concept by optimizing camera trajectories and generating sparse maps with points, lines, and planes, alongside dense surfel-based reconstructions. Similarly, [60] incorporates Manhattan frame re-identification to build robust rotational constraints, which are tightly integrated into a bundle adjustment framework. Planar-SLAM [61] focuses on planar mesh reconstruction by utilizing line and sparse point features, while [62] introduces drift-free rotation estimation through the use of Gaussian spheres. Moreover, [63] refines spatial constraints through advanced bundle adjustment leveraging structure constraints. [64] integrates low-cost LiDAR in structured scenes for better view coverage.
Despite these advancements, most existing systems leveraging structure or planar constraints rely on sparse map representations, such as point clouds or simple planar surfels, which are insufficient for reconstructing fine-grained maps with detailed textures. Additionally, the sparse map leads to discontinuities in the reconstructed map, making surface extraction and subsequent optimization difficult. To bridge this gap, we incorporate line features into the neural dense system, enhancing its tracking capabilities in indoor environments and enabling structured surface optimization.
@@ II-B SLAM with Structure Optimization
Figure 2: The two-phase pipeline illustration of the proposed MG-SLAM. The upper section visualizes the parallel processes of the tracking and mapping systems. The lower section presents the post-optimization of scene interpolation and mesh extraction.
III Method
Figure 2 illustrates the pipeline of our proposed method. Under the constraints of the Manhattan World, MG-SLAM introduces line segments and structured surfaces to enhance camera pose estimation and map reconstruction. Section III-B details the tracking mechanism that utilizes both point and line features. We utilize a specific strategy for fusing line segments to ensure reliable identification of line features. Section III-C discusses the Gaussian representation, including a specialized loss term dedicated to the reconstruction of line segments. Section III-D describes the completion and refinement of the scene, grounded in the assumption of structured surfaces. Section III-E describes the mesh generation utilizing regularization losses.
III-A Notation
@@
III-B Tracking
We utilize the backprojection of point features and line segments extracted from 2D images into 3D space in parallel for optimizing the camera pose, based on PLVS [13]. Moreover, we propose strategies for line segment fusion and suppression, designed to merge shorter segments into longer ones and eliminate unstable lines. This method is grounded in the understanding that longer line features tend to offer greater reliability and robustness throughout the tracking process.
III-B1 Point Reprojection Error
@@
III-B2 Line Reprojection and Backprojection Error
@@
$\mathcal{L}_{\rm 2D}^{l}$ $=\begin{bmatrix}\left\|d_{k}\cdot\pi(R_{k}\mathcal{X}_{p}^{w}+t)\right\|\\ \left\|d_{k}\cdot\pi(R_{k}\mathcal{X}_{q}^{w}+t)\right\|\end{bmatrix}_{\Sigma_{\rm 2D}}\in\mathbb{R}^{2}~{},$ (2)
$d_{k}$ $=\frac{\bar{\mathcal{P}_{p}}\times\bar{\mathcal{P}_{q}}}{\left\|\bar{\mathcal{P}_{p}}\times\bar{\mathcal{P}_{q}}\right\|}|_{k}$ (3)
@@
which considers the cross-product of the differences between the point’s position and the back-projected positions of the line’s endpoints, normalized by the distance between these endpoints. Additionally, the distance between the world domain point $\mathcal{X}^{w}_{i}$ and its corresponding back-projected point in the image plane $\mathcal{P}_{i}\in\mathbb{R}^{2}$ is determined by:
@@
III-B3 Line Segment Fusion
The stability of tracking can be notably enhanced by the presence of elongated line segments. However, the extraction process, particularly through EDLines, often yields fragmented segments as a consequence of various disturbances like image noise. Our method integrates the following key steps: (1) Fusing line segments that are directionally aligned within a one-degree angle difference, ensuring they follow the same path. (2) Ensuring these segments are close—within a 10-pixel distance of each other’s nearest endpoints—yet do not overlap, preserving their distinctness while allowing for precise merging. (3) Verifying the vertical distance between the corresponding endpoints of one segment to the entirety of another is less than certain pixel threshold, maintaining geometric consistency. Only segments fulfilling all three criteria are merged, producing longer and more reliable lines. Furthermore, we filter out segments that fall below a predefined length threshold, relative to the image size, to improve the system’s tracking robustness.
III-B4 Full Bundle Adjustment Error
The overall objective function for the full bundle adjustment is given by:
$\mathcal{L}_{\rm BA}=\sum_{\mathcal{K}}\sum_{\mathcal{U}}\rho\left(\left\|\mathcal{L}_{\rm 2D}^{p}\right\|_{\Sigma_{\mathcal{P}}}\right)$ $+\sum_{\mathcal{V}}\rho\left(\left\|\mathcal{L}_{\rm 2D}^{l}\right\|_{\Sigma_{\rm 2D}}\right)$
$+\sum_{\mathcal{V}}\rho\left(\left\|\mathcal{L}_{\rm 3D}^{l}\right\|_{\Sigma_{\rm 3D}}\right)$ (7)
Here, $\mathcal{K}$ represents the set of chosen keyframes, while $\mathcal{U}$ and $\mathcal{V}$ denote the sets of points and lines extracted within the current frame. The Huber cost function $\rho(\left\|e\right\|)_{\Sigma^{-1}})=e^{T}\Sigma^{-1}e$ is applied to mitigate the influence of outliers. The optimization process utilizes the Levenberg-Marquardt method [67, 68], which solves the augmented normal equations by iteratively updating the parameters $\Theta$ as:
$(A_{\epsilon}^{T}\Sigma_{\epsilon}^{-1}A_{\epsilon}+\Lambda\mathbb{I})\Delta\Theta=-A_{\epsilon}^{T}\Sigma_{\epsilon}^{-1}\epsilon$ (8)
III-C Mapping
The map representation is based on 3D Gaussians primitives $\mathcal{G}=\alpha\mathcal{N}(\mu^{w},\Sigma^{w})$, where $\alpha\in[0,1]$ is the opacity, $\mu^{w}$ and $\Sigma^{w}$ are mean and covariance in world coordinate. Each $\mathcal{G}_{i}$ is associated with the color of appearance and semantic feature $f_{i}=\{c_{i}\in\mathbb{R}^{3},s_{i}\in\mathbb{R}^{3}\}$. Semantic segmentation is utilized to identify the surface such as floors for structure optimization and surface extraction, explained in Section III-D.
III-C1 Scene Representation